OpenStudy (anonymous):

Equivalent Ratios x y 3 6 ? 10 7 14 What number is missing from the table of equivalent ratios? A) 5 B) 7 C) 13 D) 20

OpenStudy (anonymous):

@blake57roger

OpenStudy (anonymous):

Hi! I can help! Is that okay?

OpenStudy (anonymous):

yes sure @AMYCARTER

OpenStudy (anonymous):

it would be a because they r all being multiplied by 2 and 5*2=10 am i right?

OpenStudy (anonymous):

huh @AMYCARTER

OpenStudy (anonymous):

I'm not giving you the answer @Kyliejones. I am guiding you to the answer. Now since it is obvious the other ratios equal out to the same number 2, you find the number divided by 10 that equals 2. The equation you use is \[\frac{ y }{ x }\] So as you suggested let's do \[\frac{ 10 }{ 5 }\]. What does that equal to @Kyliejones?

OpenStudy (anonymous):

All right then I see you just wanted the answer @Kyliejones....

OpenStudy (anonymous):

Hun go change your fake picture @amycarter